# Ultralight Black Holes as Sources of High-Energy Particles

### v3

@@ Abstract
The memory burden effect, stating that the amount of information stored within a system contributes to its stabilization, is particularly significant in systems with a high capacity for information storage such as black holes (BHs). In these systems, the evaporation process is halted, at the latest, after approximately half of the BH’s initial mass has been radiated away. Consequently, light primordial BHs (PBHs) of mass $m_{\rm PBH}\lesssim 10^{15}$ g, which are expected to have fully evaporated by present time, may remain viable candidates for dark matter (DM). In this scenario, we demonstrate that their mergers would continue to occur today, leading to the formation of “young” BHs that resume evaporating, producing ultrahigh-energy cosmic rays detectable by current experiments. The emission spectrum would be thermal in all Standard Model particle species, offering a clear and distinguishable signature. Current measurements of the isotropic neutrino flux at Earth are in tension with light PBHs as DM candidates within the mass range $7\times 10^{3}\lesssim m_{\rm PBH}/{\rm g}\lesssim 4\times 10^{8}$. Contrarily to existing bounds, our results depend uniquely on the mass of the remnant and not on the specific model building details of the stabilized phase. We also discuss the potential for refining these constraints through gamma-ray and cosmic-ray observations, as well as gravitational wave detections.
@@
Introduction — Black holes (BHs), predicted by general relativity, are not only fascinating astrophysical objects but also intriguing sources of particle emission. A BH of mass $m_{\rm BH}$ emits high-energy particles lighter than its Hawking temperature, $T_{H}=(8\pi Gm_{\rm BH})^{-1}$ Hawking (1975). In the semiclassical regime where quantum gravity effects are neglected throughout their evaporation, the mass loss rate for an emitting BH follows $\dot{m}_{\rm BH}\propto m_{\rm BH}^{-2}$. The Hawking radiation has been searched for in terms of the BH emission into photons Carr et al. (2016); Boudaud and Cirelli (2019); DeRocco and Graham (2019); Laha (2019); Laha et al. (2020); Chan and Lee (2020), neutrinos Dasgupta et al. (2020); Bernal et al. (2022a), gravitons Hooper et al. (2020), and exotic particles Hooper et al. (2019); Schiavone et al. (2021); Auffinger (2023); Baker and Thamm (2023).
@@
However, it has been suggested that the information contained within a BH affects its decay, slowing down the evaporation process through a “memory burden” effect Dvali (2018a), which becomes significant after a portion of its mass has evaporated Dvali et al. (2020). The timescale that signals the breakdown of the semiclassical description, which is also the latest the memory burden effect sets in, reads
$\tau_{\rm sc}\simeq(1-\epsilon^{3})\,\tau_{\rm evap}\,,$ (1)
@@
where $\tau_{\rm evap}$ is the BH evaporation time in the semiclassical limit and the prefactor accounts for a fraction $1-\epsilon$ of the BH mass being lost in the process. Here, we fix $\epsilon=1/2$. The subsequent dynamics of the memory-burdened BH for times $t\gtrsim\tau_{\rm sc}$ remain uncertain due to the influence of quantum gravity effects Dvali et al. (2020). One possible outcome is that the configuration becomes unstable and eventually breaks down. Alternatively, evaporation could persist at a reduced rate, controlled by powers of the gravitational coupling, $\alpha_{\rm gr}\simeq S^{-1}$. In what follows, we assume the former scenario as the likely outcome indicated by theoretical analyses Dvali et al. (2020, 2024a).
@@
When quantum gravity effects are taken into account, the BH evaporation rate is slowed down and the BH lifetime is extended as Dvali et al. (2020)
$\tau\simeq\tau_{\rm sc}\,S^{1+k}\,,$ (2)
@@
where $S=4\pi Gm_{\rm BH}^{2}$ represents the BH entropy. Theoretical models such as the holographic description of BHs Dvali (2018b) and the microscopic description in terms of marginally bound gravitons—the so-called BH $N$-portrait Dvali and Gomez (2013a, 2014a, 2014b)- independently suggest that $k=1$ Dvali et al. (2024a), while numerical analyses give $k\geq 1$ Dvali et al. (2024a). In what follows, we will derive bounds independent of this quantity.
@@
The memory burden effect introduces a new mass window for PBHs as DM candidates below $10^{15}$ g, with BHs of mass around $10^{8}$ g shown to be viable while satisfying all known constraints Dvali et al. (2020). For $k=1$, Eq. (2) predicts that memory-burdened BHs as light as $3\times 10^{3}\rm g$ can be DM candidates. A viable mechanism for their production within this mass range $3\times 10^{4}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$, here dubbed the ultralight BH window, was introduced in Ref. Dvali et al. (2021). A more precise mass range for memory-burdened PBHs constituting DM has been recently derived Alexandre et al. (2024); Thoss et al. (2024). Other consequences of memory-burdened BHs have also been recently explored Balaji et al. (2024); Haque et al. (2024); Riotto and Silk (2024); Barman et al. (2024a, b); Bhaumik et al. (2024); Kohri et al. (2024); Jiang et al. (2024); Chianese et al. (2024); Athron et al. (2024).
In this Letter, we discuss the observational consequences of memory-burdened PBHs as DM. We argue that these PBHs, locked in binary configurations, merge to form new BHs of similar mass. Upon merging, the new object is a young BH, no longer subject to the memory burden effect. As such, the freshly-formed BH starts emitting again for a duration $\tau_{\rm sc}$, at much lower redshifts than their progenitors, thus leading to novel signatures that are unique to the memory burden phenomenon.
@@
Memory burden effect — Saturated configurations offer valuable insights into objects with high information-storage capacity, revealing connections between gravitational and non-gravitational systems. Notably, “saturons”, configurations of maximal entropy constrained by the unitarity bound Dvali (2021a), exemplify these objects. BHs are a prime example of saturated configurations, as they obey the entropy-area law Bekenstein (1973). Saturated configurations also arise in non-gravitational theories, such as renormalizable field theories Dvali (2021a, b, c). Saturated configurations not only follow an area law, but also share several properties with BHs, such as thermal-like evaporation Dvali and Sakhelashvili (2022); Dvali et al. (2022a), the presence of an information horizon in the semiclassical regime Dvali (2021b, c, a, d); Dvali and Sakhelashvili (2022); Dvali et al. (2022a), a timescale for information retrieval corresponding to the half-time of evaporation Dvali (2021a); Dvali and Sakhelashvili (2022); Dvali et al. (2022a), and a maximal spin bound that cannot exceed their entropy, similar to the extremality condition for spinning BHs Dvali et al. (2022b, 2024b). Many features traditionally associated with BHs may arise from the broader behavior of saturated configurations, rather than being unique to gravity.
This framework provides new approaches for studying BHs from both theoretical and phenomenological perspectives. By analyzing the non-gravitational counterparts of BHs, we can use the calculability of renormalizable theories to better understand their enigmatic properties. Extending the study of saturated configurations to BHs allows us to predict new phenomenological signatures. For example, rotating saturons provide a microscopic description of maximally spinning BHs as graviton condensates with vorticity Dvali et al. (2022b), offering a topological explanation for the absence of Hawking radiation. This could lead to observable deviations in the gravitational radiation from BH mergers Dvali et al. (2024b).
Furthermore, the memory burden effect, studied in calculable saturated configurations, is expected to manifest in BHs as well Dvali et al. (2020, 2024a). The memory stored in a BH can be represented as $M$ qubits, corresponding to the occupation number in the flavor space of Goldstone bosons arising from broken symmetries Dvali (2021a, b, c). The number of degenerate states is $n_{\rm st}=2^{M}$, leading to an entropy $S=\log n_{\rm st}\simeq M$. As BHs decay, the increasing mass gap of the Goldstone bosons halts evaporation by the half-mass decay time. [^1] The memory burden effect is an explicit realization of the so-called quantum-breaking of classical configurations Dvali et al. (2013), discussed within the framework of the BH $N$-portrait Dvali and Gomez (2013b, 2014b), which questions the validity of the classical description of BHs and finds that it is violated around $\tau_{\rm sc}$. This can be modeled using the probability distribution for memory-occupied modes Dvali et al. (2024a),
@@
where $N_{\rm G}$ represents the order parameter setting the timescale of quantum backreaction Dvali et al. (2024a). The probability distribution in (3) leads to a natural spread in the mass distribution of PBHs Dvali et al. (2024a), offering a unique signature and further insights into BH evolution and mergers. A careful dynamical analysis reveals that an evaporating BH asymptotically evolves into a memory-burdened BH with around half of its initial mass Dvali et al. (2024a). While it is possible for a BH with empty memory to exist, allowing for its complete evaporation without memory burden, such cases are exponentially rare, as indicated by the probability distribution in Eq. (3).
@@
As suggested by the study of mergers Dvali et al. (2024b) of memory-burdened configurations Dvali et al. (2024a), only the merging point is affected by the memory burden, potentially emitting additional radiation, while the merged object is unaffected. Mapped to the gravitational counterpart, these results imply that the memory burden of the two merging BHs leads to significant deviations in the gravitational wave signal only at ringdown, when a new classical BH forms with a memory pattern governed by Eq. (3). Besides, the newly formed BH has an area, and thus an entropy, approximately four times larger than the combined areas of the two initial BHs, providing considerable memory to store all the non-emitted information. Therefore, post-merger the BH emits semiclassically for a time $\tau_{\rm sc}$ before stabilizing again. [^2] The merger of extremal ultralight PBHs was discussed in Ref. Bai and Orlofsky (2020) where it was argued that Hawking emission is reactivated. Indeed, the memory burden effect shares similarities with BHs that are semiclassically stabilized by extremality Dvali et al. (2024a). This leads to important phenomenological consequences, which are discussed next.
@@
The cosmic ray flux from evaporating PBHs — The present fraction of memory-burdened PBHs formed at temperature $T_{f}$ is related to their primordial abundance $\beta$ through the conservation of the PBH comoving number, as
@@
The differential rate of PBH binary mergers describes a distribution of PBHs that decoupled from the Hubble flow before matter-radiation equality Ali-Haïmoud et al. (2017); Raidal et al. (2019); Liu et al. (2019). For a monochromatic mass function, the most conservative merger rate is Vaskonen and Veermäe (2020); Hütsi et al. (2021); Jedamzik (2020); Young and Hamers (2020); Jedamzik (2021)
$R_{\rm PBH}(t)=\frac{4.9\times 10^{16}}{\rm kpc^{3}\,yr}\,f_{\rm PBH}^{\frac{53}{37}}\,\left(\frac{t_{0}}{t}\right)^{\frac{34}{37}}\,\left(\frac{m_{\rm PBH}}{\rm g}\right)^{-\frac{32}{37}}S\,,$ (5)
where the suppression factor $S$ is the product of two components: a term $S_{1}\approx 0.24$ which accounts for the interactions between the binary system and the surrounding DM inhomogeneities, as well as neighboring BHs Hütsi et al. (2021), and $S_{2}(x)\approx\min\left[1,9.6\times 10^{-3}x^{-0.65}\exp\left(0.03\ln^{2}x\right)\right]$ parameterising the suppression caused by BHs absorbed by collapsed PBH clusters. Here, $x\equiv\left(t/t_{0}\right)^{0.44}f_{\mathrm{PBH}}$. The function $S_{2}(x)$ attains a minimum at $10^{-2}$, consistently with numerical simulations Inman and Ali-Haïmoud (2019); Franciolini et al. (2022). For light PBHs, dynamical captures induced by gravitational waves (GWs) as well as late-time dynamical capture are negligible Franciolini et al. (2022). However, local non-Gaussianity in primordial curvature perturbations can cluster PBHs at formation, potentially enhancing the merger rate by up to $\mathcal{O}(10^{7})$ Raidal et al. (2017); De Luca et al. (2021); Franciolini et al. (2022). Although some level of clustering is expected from various mechanisms at the formation of the PBH binary, we consider the most conservative rate expressed by Eq. (5) as the fiducial model.
@@
where $\delta(r)\equiv\rho_{\rm DM,gal}(r)/\rho_{\rm DM}$ describes the Galactic enhancement in the DM density distribution Pujolas et al. (2021), which is here modeled as a Navarro-Frenk-White (NFW) profile Navarro et al. (1997), $\rho_{\rm DM,gal}(r)=\rho_{s}(r_{s}/r)(1+r/r_{s})^{-2}$, where we fix $r_{s}=15.6$ kpc and $\rho_{s}$ is taken so that the DM density in the vicinity of the Solar System is $\rho_{\odot}=0.45{\rm\,GeV\,cm^{-3}}$, while $r_{\rm vir}$ is defined as $\delta(r_{\rm vir})=200$. The function $r(s,\theta)$ accounts for the line-of-sight distance $s$ of the merger from Earth, with the integration over $s$ extending up to the value for which $r=r_{\rm vir}$. Averaging the differential flux over the angular direction $\theta$ and integrating in the energy bins yields the integral flux $\Phi_{i}^{\rm gal}(E)$. The extra-galactic (eg) contribution to the differential flux of the species $i$ is
@@
where ${\rm d}t/{\rm d}z$ is the cosmological line element, $E^{\prime}=E\,(1+z)$ accounts for the redshift of the energy $E^{\prime}$ at emission and $z_{f}$ is the redshift of PBHs at formation. The opacity $\eta_{i}$ describes the attenuation between $z$ and today at energy $E$.

### v4

@@ Abstract
The memory burden effect, the idea that the amount of information stored within a system contributes to its stabilization, is particularly relevant for systems with a large information storage capacity, such as black holes. In these objects, the evaporation process halts, at the latest, once approximately half of the initial mass has been radiated away. As a result, light primordial black holes (PBHs) with mass $m_{\rm PBH}\lesssim 10^{15}\,\mathrm{g}$, which are traditionally assumed to have fully evaporated by the present time, may instead survive and constitute viable dark matter candidates. Ongoing mergers of such PBHs would give rise to “young” black holes that resume their evaporation, emitting ultrahigh-energy particles potentially detectable by current experiments. The resulting emission spectrum would be thermal across all Standard Model particle species, offering a clear and distinctive signature. We demonstrate that, if the memory burden effect activates after PBHs have lost around half of their initial mass, current measurements of the neutrino flux at Earth place strong constraints on such dark matter candidates for $m_{\rm PBH}\lesssim 10^{9}\,\mathrm{g}$. This suggests that the memory burden must set in at earlier stages of evaporation. Unlike existing bounds, our results depend solely on the mass of the remnant, and not on model-dependent details of the stabilized phase. We also discuss the potential for refining these constraints through observations of gamma rays, cosmic rays, and gravitational waves.
I Introduction
@@
Black holes (BHs), predicted by general relativity, are not only fascinating astrophysical objects but also intriguing sources of particle emission. A BH of mass $m_{\rm BH}$ emits high-energy particles lighter than its Hawking temperature, $T_{\rm H}=(8\pi Gm_{\rm BH})^{-1}$ Hawking (1975). In the semiclassical regime where quantum gravity effects are neglected throughout their evaporation, the mass loss rate for an emitting BH follows $\dot{m}_{\rm BH}\propto m_{\rm BH}^{-2}$. The Hawking radiation has been searched for in terms of the BH emission into photons Carr et al. (2016); Boudaud and Cirelli (2019); DeRocco and Graham (2019); Laha (2019); Laha et al. (2020); Chan and Lee (2020), neutrinos Dasgupta et al. (2020); Bernal et al. (2022a), gravitons Hooper et al. (2020), and exotic particles Hooper et al. (2019); Schiavone et al. (2021); Auffinger (2023); Baker and Thamm (2023).
@@
However, it has been suggested that the information contained within a BH affects its decay, slowing down the evaporation process through a “memory burden” effect Dvali (2018a); Dvali et al. (2020, 2024a), which becomes significant after a portion of its mass has evaporated Dvali et al. (2020, 2024a). The timescale that signals the breakdown of the semiclassical description and the set in of memory burden is
$\tau_{\rm sc}\simeq q\,\tau_{\rm evap}\,,$ (1)
@@
where $\tau_{\rm evap}$ is the BH evaporation time according to the semiclassical rate and the prefactor accounts for the fraction $q\doteq\Delta m_{\rm PBH}/m_{\rm PBH}$ of the BH mass being lost during the semiclassical emission phase. Here, we fix $q=1/2$, corresponding to the latest time that memory burden may kick in. The subsequent dynamics of the memory-burdened BH for times $t\gtrsim\tau_{\rm sc}$ remain uncertain due to the influence of quantum gravity effects Dvali et al. (2020, 2024a). One possible outcome is that the configuration becomes unstable and eventually breaks down. Alternatively, evaporation could persist at a reduced rate, controlled by powers of the gravitational coupling, $\alpha_{\rm gr}\simeq S^{-1}$. In what follows, we assume the former scenario as the likely outcome which is further indicated by theoretical analyses Dvali et al. (2020, 2024a).
@@
When quantum effects are taken into account, the BH evaporation rate is slowed down and the BH lifetime is extended as Dvali et al. (2020)
$\tau\simeq\tau_{\rm evap}\,S^{k}\,,$ (2)
@@
where $S=4\pi Gm_{\rm BH}^{2}$ represents the BH entropy. Theoretical models such as the holographic description of BHs Dvali (2018b) and the microscopic description in terms of marginally bound gravitons—the so-called BH $N$-portrait Dvali and Gomez (2013a, 2014a, 2014b)- independently suggest that $k=2$ Dvali et al. (2020, 2024a) (consistency requires $k\geq 1$). In what follows, we will derive bounds independent of this quantity.
@@
The memory burden effect introduces a new mass window for PBHs as DM candidates below $10^{15}$ g, with BHs of mass around $10^{8}$ g shown to be viable while satisfying all known constraints Dvali et al. (2020). For $k=2$, Eq. (2) implies that memory-burdened BHs as light as $\gtrsim 10^{3}\rm g$ can be DM candidates. A viable mechanism for their production within this mass range $3\times 10^{4}\lesssim m_{\rm PBH}/{\rm g}\lesssim 10^{9}$, here dubbed the ultralight BH window, was introduced in Ref. Dvali et al. (2021). [^1] See also Refs. Carr et al. (1994); Profumo (2025) for Planck-scale relics. A more precise mass range for memory-burdened PBHs constituting DM has been recently derived Alexandre et al. (2024); Thoss et al. (2024). Other consequences of memory-burdened BHs have also been recently explored Balaji et al. (2024); Haque et al. (2024); Riotto and Silk (2024); Barman et al. (2024a, b); Bhaumik et al. (2024); Kohri et al. (2025); Jiang et al. (2024); Chianese et al. (2025); Athron et al. (2025); Dvali et al. (2025).
In this paper, we discuss the observational consequences of memory-burdened PBHs as DM. We argue that these PBHs, locked in binary configurations, merge to form new BHs of similar mass. Upon merging, the new object is a young BH, no longer subject to the memory burden effect. As such, the freshly-formed BH starts emitting again for a duration $\tau_{\rm sc}$, at much lower redshifts than their progenitors, thus leading to novel signatures that are unique to the memory burden phenomenon.
@@ II.1 Memory burden effect
Saturated configurations offer valuable insights into objects with high information-storage capacity, revealing connections between gravitational and non-gravitational systems. Notably, “saturons”, configurations of maximal entropy constrained by the unitarity bound Dvali (2021a), exemplify these objects. BHs are a prime example of saturated configurations, as they obey the entropy-area law Bekenstein (1973). Saturons also arise in non-gravitational theories, such as renormalizable field theories Dvali (2021a, b, c). These objects not only follow an area law, but also share several properties with BHs, such as thermal-like evaporation Dvali and Sakhelashvili (2022); Dvali et al. (2022a), the presence of an information horizon in the semiclassical regime Dvali (2021b, c, a, d); Dvali and Sakhelashvili (2022); Dvali et al. (2022a), a timescale for information retrieval corresponding to the half-time of evaporation Dvali (2021a); Dvali and Sakhelashvili (2022); Dvali et al. (2022a), and a maximal spin bound that cannot exceed their entropy, similar to the extremality condition for spinning BHs Dvali et al. (2022b, 2024b). As a consequence, many features traditionally associated with BHs may arise from the broader behavior of saturated configurations, rather than being unique to gravity.
This framework provides new approaches for studying BHs from both theoretical and phenomenological perspectives. By analyzing the non-gravitational counterparts of BHs, we can use the calculability of renormalizable theories to better understand their enigmatic properties. Moreover, the study of this class of objects, allows for the extension of novel properties to BHs by universality.
The memory burden phenomenon is a necessary feature of all saturons, including BHs. This stems from the requirement of generating a large microstate degeneracy to account for the entropy-area law. Achieving this requires a large number of modes with zero energy gap or “gapless”, localized on the support of the BH. Outside the object, these same degrees of freedom become gapped. As a result, for the BH to release its memory, an energy barrier needs to be overcome. Hawking radiation cannot do this due to its softness. The very same energy barrier is the one the BH eventually has to face as Hawking evaporation reduces its mass. We shall not rederive the memory burden mechanism here, and instead refer the reader to the relevant literature Dvali (2021a, b, c). Our focus is on features relevant for the merger process.
The memory stored in a BH can be represented as $M$ qubits, which, due to their gaplessness, leads to a number of degenerate microstates $n_{\rm st}=2^{M}$, and therefore an entropy $S=\log n_{\rm st}\simeq M$. The amount of memory stored within the BH, corresponding to the total number of occupied memory modes $N_{\rm G}$, is determined statistically, according to the probability distribution Dvali et al. (2024a),
@@
where $N_{\rm G}$ is effectively the order parameter responsible for the quantum backreaction Dvali et al. (2024a). On average, $N_{\rm G}\simeq S/2$. As BH decays, the increasing mass gap of the memory modes eventually halts the evaporation process. A careful dynamical analysis reveals that this transition is universally determined solely by the critical exponent, and can occur within a timescale $\tau_{\rm evap}/\sqrt{S}\lesssim\tau_{\rm SC}\lesssim\tau_{\rm evap}/2$ Dvali et al. (2020, 2024a). [^2] The memory burden effect is an explicit realization of the so-called quantum-breaking of classical configurations Dvali et al. (2013), discussed within the framework of the BH $N$-portrait Dvali and Gomez (2013b, 2014b), which questions the validity of the classical description of BHs and finds that it is violated around $\tau_{\rm sc}$.
@@
The probability distribution in (3) leads to a natural spread in the mass distribution of PBHs Dvali et al. (2024a), offering a unique signature and further insights into BH evolution and mergers. While it is possible for a BH with empty memory to exist, allowing for its complete evaporation without memory burden, such cases are exponentially rare, as indicated by the probability distribution in Eq. (3).
When two memory-burdened BHs merge, we expect a “young” BH to form, due to the large phase space volume occupied by these configurations. In fact, semiclassical BHs are realized when the memory modes are gapless and cost zero energy. At this point, the newly formed object possesses a large microstate degeneracy, and all such configurations need to be included in the amplitude, effectively saturating it Dvali (2021a). This saturation does not occur if one considers memory burdened BHs as the outcome, since BHs with different macroscopic quantum hairs have different energies and therefore do not saturate the phase space Dvali et al. (2024a). Equivalently stated, the formation of an initially classical BH is largely insensitive to the initial pre-merger configuration, due to the large entropy involved.
A natural question is whether memory burden could dynamically induce significant deviations in the mass of the final object, compared to the initial configuration. The answer appears to be negative. In fact, as suggested by studies of mergers Dvali et al. (2024b) of memory-burdened configurations Dvali et al. (2024a), only the merging point is altered by the memory modes, and can lead to the radiation of a small mass fraction. This, however, may result in a final BH with a spin configuration that significantly differs compared to the case without memory. Mapped to the gravitational counterparts, these results suggest that the memory burden of the two merging BHs may source significant deviations in the gravitational wave signal during the ringdown phase, with the resulting BH showing a different angular momentum than classically expected. Since Hawking radiation after the merger is sensitive to the rotation of the object, which is also determined by the initial unknown spin of the two progenitors, we restrict our analysis to non-spinning BHs. The inclusion of rotational effects would strengthen our results.
@@
Although the merger dynamics can also lead to a partial release of the information, the newly formed BH has an area, and thus an entropy, approximately four times larger than the combined areas of the two initial BHs, providing considerable memory-storage capacity for all the initial information. Therefore, we expect that a new classical BH forms with a memory pattern governed by Eq. (3). As a consequence, post-merger the BH emits semiclassically for a time $\tau_{\rm sc}$ before stabilizing again. [^3] The merger of extremal ultralight PBHs was discussed in Ref. Bai and Orlofsky (2020) where it was argued that Hawking emission is reactivated. Indeed, memory-burdened PBHs may share some phenomenological similarities with BHs that are semiclassically stabilized by extremality Dvali et al. (2024a). This leads to important phenomenological consequences, which are discussed next.
II.2 High-energy particles from evaporating PBHs
@@
The present fraction of memory-burdened PBHs formed at temperature $T_{f}$ is related to their primordial abundance $\beta$ through the conservation of the PBH comoving number, as
@@
The differential rate of PBH binary mergers describes a distribution of PBHs that decoupled from the Hubble flow before matter-radiation equality Ali-Ha¨ımoud et al. (2017); Raidal et al. (2019); Liu et al. (2019). For a monochromatic mass function, the most conservative merger rate is Vaskonen and Veermäe (2020); Hütsi et al. (2021); Jedamzik (2020); Young and Hamers (2020); Jedamzik (2021)
$R_{\rm PBH}(t)\simeq\frac{5.7\times 10^{-66}}{\rm cm^{3}\,s}f_{\rm PBH}^{\frac{53}{37}}\left(\frac{t_{0}}{t}\right)^{\frac{34}{37}}\left(\frac{2m_{\rm PBH}}{10^{10}\rm g}\right)^{-\frac{32}{37}}S\,,$ (5)
where the suppression factor $S$ is the product of two components: a term $S_{1}\approx 0.24$ which accounts for the interactions between the binary system and the surrounding DM inhomogeneities, as well as neighboring BHs Hütsi et al. (2021), and $S_{2}(x)\approx\min\left[1,9.6\times 10^{-3}x^{-0.65}\exp\left(0.03\ln^{2}x\right)\right]$ parameterising the suppression caused by BHs absorbed by collapsed PBH clusters. Here, $x\equiv\left(t/t_{0}\right)^{0.44}f_{\mathrm{PBH}}$. The function $S_{2}(x)$ attains a minimum at $10^{-2}$, consistently with numerical simulations Inman and Ali-Ha¨ımoud (2019); Franciolini et al. (2022). For light PBHs, dynamical captures induced by gravitational waves (GWs) as well as late-time dynamical capture are negligible Franciolini et al. (2022). However, local non-Gaussianity in primordial curvature perturbations can cluster PBHs at formation, potentially enhancing the merger rate by up to $\mathcal{O}(10^{7})$ Raidal et al. (2017); De Luca et al. (2021); Franciolini et al. (2022). Although some level of clustering is expected from various mechanisms at the formation of the PBH binary, we consider the most conservative rate expressed by Eq. (5) as the fiducial model.
[…]
@@
where $\delta(r)\equiv\rho_{\rm DM,gal}(r)/\rho_{\rm DM}$ describes the Galactic enhancement in the DM density distribution Pujolas et al. (2021), which is here modeled as a Navarro-Frenk-White (NFW) profile Navarro et al. (1997), $\rho_{\rm DM,gal}(r)=\rho_{s}(r_{s}/r)(1+r/r_{s})^{-2}$, where we fix $r_{s}=15.6$ kpc and $\rho_{s}$ is taken so that the DM density in the vicinity of the Solar System is $\rho_{\odot}=0.45{\rm\,GeV\,cm^{-3}}$, while $r_{\rm vir}$ is defined as $\delta(r_{\rm vir})=200$. The function $r(s,\theta)$ accounts for the line-of-sight distance $s$ of the merger from Earth, with the integration over $s$ extending up to the value for which $r=r_{\rm vir}$. Averaging the differential flux over the angular direction $\theta$ and integrating in the energy bins yields the integral flux $\Phi_{i}^{\rm gal}(E)$. The extra-galactic (eg) contribution to the differential flux of the species $i$ is [^5] We thank Paolo Panci for noticing a typo in the original formula.
@@
where ${\rm d}t/{\rm d}z$ is the cosmological line element, $E^{\prime}=E\,(1+z)$ accounts for the redshift of the energy $E^{\prime}$ at emission and $z_{f}$ is the redshift of PBHs at formation. The opacity $\eta_{i}$ describes the attenuation between $z$ and today at energy $E$. More details can be found in Dondarini et al. (2025).
III Results